Question 5 2 pts Suppose the wavelength in the formula developed by Balmer was 410 nm. What is the value of n? O 2 O 3 0 4 0 5 0 6
Answer #1

1/wavelength = RH [1/2^2 - 1/n^2]

where RH = 1.0967 x 10^7 m^-1

1/(410x 10^-9) = (1.0967 x 10^7) [ 1/4 - 1/n^2]

1/4 - 1/n^2 = 0.2224

1/n^2 = 0.0276

n = 6